what are the solutions of sin(4x) + sin(x) = 0 between 0 and pi
Found 2 solutions by addingup, ikleyn:
Answer by addingup(3677)

sin(4x) + sin(x) = 0
Rewrite the equation:
4x = (sin^−1)(sin(x)). The right hand side is either x or π−x, or either of these plus any integer multiple of 2π. So, in the following equation find all values of n that give results between 0 and π:
4x = x + 2nπ
or
4x=π−x + 2nπ
~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~
example:
x = (2nπ)/3 or x = ((2n+1)π)/5

Answer by ikleyn(53937)

.

            What I see in the post by @addingup,  is not a correct solution.

            It is even not a  Math  -  it is far away from any  Math standard.

            You can  (actually,  you  MUST)  ignore it  -  it does not deserve any discussion.


For a correct solution see my post under the link

https://www.algebra.com/algebra/homework/Trigonometry-basics/Trigonometry-basics.faq.question.1127695.html

https://www.algebra.com/algebra/homework/Trigonometry-basics/Trigonometry-basics.faq.question.1127695.html